the sequence 24, a, b, c, 60 is arithmetic. What is the value of a+b+c?
Answer by John10(297)

Find a, b,c
the arithmetic sequence is an = a1 + (n - 1)d
we have a1 and a5 which are 24 and 60 respectively
a5 = a1 + (5 -1)d
60 = 24 + 4d
4d = 60 - 24
4d = 36
d = 36/4
d = 9
Thus a = a1 + d = 24 + 9 = 33
b = a1 + 2d = 24 + 2(9) = 24 + 18 = 42
c = a1 + 3d = 24 + 3(9) = 24 + 27 = 51
Then their sum is a + b + c = 33 + 42 + 51 = 126
John10:)